A patient receives 3 liters of an electrolyte fluid containing 234 mg of sodium chloride (NaCl—m.w. 58.5), 125 mg of potassium acetate (C2 H3 KO2 —m.w. 98), and 21 mg of magnesium acetate (C4 H6 MgO4 —m.w. 142) per 100 mL. How many milliequivalents each of Na+, K+, and Mg++ does the patient receive?
Answer #1

3 liters of electrolyte fluid contains

234 mg of sodium chloride in 100ml

125 mg of potassium acetate in 100ml

21mg of magnesium acetate in 100ml

3 liters equal to 3000 ml

# 100 ml of electrolyte fluid contain 234 mg of sodium chloride

1ml of electrolyte fluid contain 234 / 100ml of NaCl

3000 ml of electrolyte fluid contain 234/ 100 × 3000

= 7020 mg

# 100 ml of electrolyte fluid contain 125 mg of potassium acetate

1ml of electrolyte fluid will have 125/ 100 mg of potassium acetate

3000 ml of electrolyte fluid contain 125/ 100 × 3000

= 3750 mg

# 100 ml of electrolyte fluid contain 21 mg of magnesium acetate

1ml of electrolyte fluid will contain 21/ 100 mg of magnesium acetate

3000ml of electrolyte fluid contain 21/ 100× 3000

= 630 mg

So, 3 liters of electrolyte fluid contains :-

Sodium chloride = 7020 mg

Potassium acetate = 3750 mg

Magnesium acetate = 630 mg

Then ,

23 mg of Na is equal to 1 mEq

1mg of Na is equal to 1/ 23mEq

7020 mg of Na= 1/ 23× 7020 = 305 .2 mEq

Sodium = 305.2mEq

39 mg of potassium is equal to 1 mEq

1 mg of potassium is equal to 1/ 39 mEq

3750mg of potassium = 1/ 39 × 3750 mEq = 96.1 mEq

Potassium = 96.1 mEq

1gm of Mg is equal to 8.12 mEq

1mg of Mg is equal to 8.12 / 1000 mEq

630 mg of Magnesium = 8.12/1000 × 630 = 5.1mEq

Magnesium = 5.1 mEq